Please help me solve the system
2x-3y+4z= -14
4x+2y-3z= -17
6x-5y+7z= -37
find x,y,z

Found 2 solutions by Alan3354, Boreal:
Answer by Alan3354(69443)

Similar problem:
=========================================================
Please help me solve the system
3x+2y-2z=14
2x-5y+3z=7
4x-3y+7z=5
================
I would eliminate x first.
Multiply the 1st eqn by 2 and the 2nd by 3, then subtract.
--> eqn in y & z
====================
Multiply the 2nd eqn by 2, subtract from the 3rd eqn.
--> another eqn in y & z
===============

Answer by Boreal(15235)

Set up the determinant matrix, which are the coefficients of the variables in a 3 x 3 matrix. Find the determinant, which is 8.
Put the column of constants in the first column of the above matrix, replacing those values with -14,-17, and -37. This will give the value for x, when divided by the 8 above.
Det (x)=-40/8=-5
Similarly, do y
Det (y)=-0
Det (z)=-8/8=-1
(-5, 0, -1) ANSWER
Check-10-4=-14
-20-(-3)=-17
-30-7=-37